- Remote participants by eZuce: None
Report of work (*:responsible person)
- Alignment, ETMY: *Enomoto, Miyo, Miyamoto, Izumi
- Tuning of ETMY damping is still going on.
- Beam alignment to both end and REFL port was confirmed the last night.
- Beam is now being drifted to down slowly at PSL by ~100urad.
- One of the DAC channel of IMMT puts half voltage. We changed the channel position in the software.
- When the bottom screws was tightened, oplev showed the shift of IMMT1 at least.
- BS: Top of the BS chamber was tightened but no big shift was not found.
- AOS: IYC and IYA started pumping down.
- ETMY: *Enomoto
- BF of tower of ETMY matrix was messed up. Polarity was changed by phase shift or something.
- CRY: Mirror reached at 18K.
Work plan (*:responsible person)
- Alignment: Enomonoto, Izumi
- IMMT coil driver will be checked.
- REFL camera has been placed and will be setup today.
- VAC: Uchiyama
- Duct connection.
- Camera: Yokozawa
- Both door will be closed today. Pump down will be start today.
- AOS: IYC will be closed and start pump-down.
- [AM] IR: PSL, Green: off.
- [PM] IR: PSL, Green: off.
- [night] IR: all, Green: off
- week of 3/12
- 3/14: meeting for pump-down check
- TMS installation at ETMY.
- Connecting ducts at the corner area.
- week of 3/19
- Vacuum test at at the corner area.
- TMS alignment using green light at ETMY.
- Beam alignment to REFL port.
- Connection of PD signal to DAQ should be implemented this week or next week.